Determine the area of a 6-inch × 10-inch rectangle by using the following units of area:
(i) A 1-inch × 1-inch square square; (ii) A 1-inch × 2-inch square square; and (iii) A right triangle
with legs 1-inch × 1-inch.

Answer by checkley77(12844)

6*10=60 in^2 is the total area.
A 1X1 square=1 in^1 so 60/1=60 squares are needed.
A 1X2= 2 in^2so 60/2=30 pieces are needed
A triangle with 1 in legs=1/2 in^2 so 60/(1/2)=60*2=120 pieces are needed.